# Methods Paper Version 19

### v1

@@ IV.2 Two-channel results
As in the single-channel case, one cannot expect a good description of the wave function if the scattering phase shifts and mixing parameter are not reproduced. We present in Figs. 20 and 21 absolute magnitudes and arguments of the wave function expansion coefficients $d_{n,ij}(E)$ respectively, at the energy of $E=0.5$ MeV where we obtain $\delta_{0}$, $\delta_{2}$, and $\epsilon$ close to exact value in calculations with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and ${}^{3}D_{1}$ as the first channel with $i=1$. We observe a fast convergence of the amplitudes even with small $N$ values.
In a wider range of energies, the wave function expansion coefficients $d_{n,ij}(E)$ obtained in calculations with $N_{\max}=12$ start deviating from the exact values because the phase shift $\delta_{0}$, $\delta_{2}$ and the mixing parameter $\epsilon$ descriptions are not accurate. However, the description of expansion coefficients $d_{n,ij}(E)$ may be improved by switching to larger $N_{\max}$ values. For example, in calculations with $N_{\max}=20$, we obtain excellent expansion coefficients $d_{n,ij}(E)$ up to an energy of 50 MeV using the set of SRFs $N=11\,\{0,1,2,3,4,5,6,7,8\}$, which provides an accurate description of scattering parameters as is seen in Fig. 19.
V ${S}$-matrix poles
We present and discuss calculations of the $S$-matrix poles within the HORSE and Efros approaches using three simple examples. In particular, we explore a resonance in $n\alpha$ scattering using the Woods–Saxon interaction with parameters fitted in Ref. WoodsSaxon1979 . Next, we consider an $S$-matrix pole associated with a bound state using as an example the deuteron bound state supported by the Daejeon16 $NN$ interaction. Finally, we analyze a narrow resonance in a model two-channel problem originally presented in Ref. Norro1980 and used as a benchmark for different numerical techniques in Refs. Moiseyev1990 ; Rakityansky2006 .
V.1 Single-channel resonance $S$-matrix pole
To describe the $n\alpha$ scattering with the Woods–Saxon interaction fitted in Ref. WoodsSaxon1979 , we use the reduced mass of $m=\frac{4}{5}m_{N}=751.136~{}{\textrm{MeV}}/{c^{2}}$, ${\hbar\Omega=30}$ MeV, and the smoothing parameter $a=5$. We investigate a resonance in the ${}^{2}P_{\frac{3}{2}}$ partial wave by locating numerically the corresponding $S$-matrix pole. Figure 22 shows the convergence with $N_{\textrm{max}}$ of the resonance energy $E_{r}$ and width $\Gamma$ obtained by the HORSE method which is equivalent to the Efros method in the case of a “complete” set of SRFs as discussed above; thus, to locate the $S$-matrix pole, one needs to solve Eq. (46) in the complex momentum plane. We see that at $N_{\textrm{max}}=10$, we obtain results very close to the exact values of $E_{r}=0.837$ MeV and $\Gamma=0.780$ MeV (see Refs. AMShirokovHORSE2016 ; BlokhintsevSSHorse ). At larger $N_{\max}$, there are small oscillations of both $E_{r}$ and $\Gamma$ around the exact values, and starting from $N_{\max}=20$, the energy and width are stable and coincide with the exact values up to at least three significant digits.
In the search for the $S$-matrix pole by the Efros method, which can use a smaller number of eigenstates, we perform the calculations with the set of the SRF eigenfunctions with increasing eigenvalues $N\,\{0,1,2,...\}$ at ${N_{\textrm{max}}=12}$ and compare with the result obtained within the HORSE method. The results are presented in Fig. 23. Both $E_{r}$ and $\Gamma$ exhibit oscillations, but note that the scale in Fig. 23 is significantly expanded as compared with that of Fig. 22. The energy and width are seen to converge with $N$ reasonably fast to the exact values and starting from $N=5$, differ from the exact values by less than 5%.
V.2 Bound state $S$-matrix pole
@@
We next search for the $S$-matrix pole in the coupled ${}^{3}SD_{1}$ partial waves associated with the deuteron bound state energy $E_{b}$ using the Daejeon16 $NN$ interaction smoothed with the parameter $a=5$ and $\hbar\Omega=30$ MeV. We first utilize the HORSE method and present in Fig. 24 the convergence of $E_{b}$ with $N_{\textrm{max}}$. The deuteron ground state energy is seen to converge much faster to the exact value than the results of the direct diagonalization of the truncated Hamiltonian $H^{tr}$. In particular, in the case of the direct diagonalization, approximately $N_{\textrm{max}}=50$ is needed to achieve the accuracy of three significant digits. In contrast, by the $S$-matrix pole location via the HORSE method, we obtain the same accuracy at ${N_{\textrm{max}}=14}$, and the accuracy of four significant digits is achieved at $N_{\max}=20$.
We obtain also a fast convergence by locating the $S$-matrix pole using the Efros method which utilizes a limited number of SRF eigenfunctions with increasing eigenvalues $N\>\{0,1,2,...\}$. The results for the ground state energy $E_{b}$ obtained with $N_{\textrm{max}}=12$ are presented in Fig. 25. $E_{b}$ exhibits jumps and instability as $N$ increases up to $N=12$; note, however, an extended scale in Fig. 25. In particular, even at small $N$ starting from $N=3$, the fluctuations of $E_{b}$ never deviate by more than 4% from the exact value.
@@
We present in Fig. 26 $\breve{d}_{n,i}(E_{b})$ of the deuteron wave function obtained by the Efros method with different $N$ using eigenfunction SRFs, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$. We see a high accuracy of expansion coefficients calculated even with small $N$.
We note that with the diagonalization of the truncated Hamiltonian matrix $H^{tr}$, we obtain the expansion coefficients $\breve{d}_{n,i}(E_{b})$ only with $n\leq\mathcal{\color[rgb]{0,0,0}N}_{i}-1$. That is, in our case of $N_{\max}=12$, we have $n\leq 6$ in the ${}^{3}S_{1}$ and $n\leq 5$ in the ${}^{3}D_{1}$ partial waves. Within the HORSE and Efros methods, we obtain coefficients $\breve{d}_{n,i}(E_{b})$ with $n$ ranging up to infinity in each channel. The coefficients $\breve{d}_{n,i}(E_{b})$ with $n>\mathcal{\color[rgb]{0,0,0}N}_{i}-1$ decrease monotonically and are responsible for the description of asymptotic tails of the bound state wave functions $A_{i}e^{-\varkappa r\!}$, where $A_{i}$ is the so-called asymptotic normalization constant in the channel $i$ and $\varkappa=\frac{1}{\hbar}\sqrt{2mE_{b}}$. The description of the asymptotic tails of bound state wave functions is important especially in weakly-bound systems such as the deuteron since they contribute to the rms radius of the system, to its quadrupole moment, to electromagnetic transitions, etc. The asymptotic normalization constants $A_{i}$ are important observables related to the residue of the $S$-matrix pole Baz which are widely used in calculations of various reactions.
V.3 Two-channel resonance
@@
$V_{ij}=\lambda_{ij}r^{2}e^{-r}+E_{i}^{\textrm{thresh.}}\delta_{ij},$ (88a)
where
$\lambda_{11}=1,\quad\lambda_{12}=\lambda_{21}=-7.5,\quad\lambda_{22}=7.5,$ (88b)
and the threshold energies in the channels 1 and 2 are
$E_{1}^{\textrm{thresh.}}=0,\quad E_{2}^{\textrm{thresh.}}=0.1.$ (88c)
The atomic units are used here.
@@
The interaction (88) has a rich spectrum containing four bound states and a number of resonances, and is therefore a good test problem for the numerical validity of a particular approach. We explore the lowest-lying narrow resonance in this system at the energy $E_{r}=4.7682$ au with the width $\Gamma=0.00142$ au.
@@
We present in Fig. 27 the resonant energy $E_{r}$ and width $\Gamma$ obtained by the HORSE method with ${\Omega=1.5}$ au, smoothing parameter $a=5$, and various $N_{\max}$. The energy $E_{r}$ is seen to converge at relatively small $N_{\max}$. In particular, $N_{\textrm{max}}=10$ is sufficient for obtaining $E_{r}$ with an accuracy of 1%. For the width $\Gamma$, approximately $N_{\textrm{max}}=50$ is needed to achieve a similar level of accuracy. Note, however, that the resonance is very narrow and the corresponding $\Gamma$ is small; thus the accuracy of 1% for the width means fixing it within a very small interval of absolute values. For both of these observables, we see a good level of stability beyond $N_{\max}=22$.
@@
The results of calculations of $E_{r}$ and $\Gamma$ within the Efros method with eigenfunction SRFs, $\Omega=1.5$ au, $N_{\max}=20$, and smoothing parameter $a=5$ at various $N$ are presented in Fig. 28 and are compared with the HORSE method result. We see that the level of accuracy within better than 0.1% of the exact result is achieved for $E_{r}$ at $N=10$. For the width $\Gamma$, a relative accuracy of better than 15% is achieved at around $N=19$. We note that both the resonance energy $E_{r}$ and its width $\Gamma$ converge with $N$ to the HORSE results which require larger $N_{\max}$ values to reproduce accurately the width. With $N=19$ we nearly exactly reproduce the HORSE width; thus one needs to perform calculations with larger $N_{\max}$ to describe the exact $\Gamma$ value with higher accuracy.
VI Conclusions
We investigated in detail possibilities of the Efros method VDEfros in describing scattering parameters, wave functions, and in locating $S$-matrix poles to obtain resonant energies and widths or accurate bound state energies and wave functions. To study the convergence and accuracy of the Efros method, we utilized various two-body systems where the exact results may be obtained by other approaches. In particular, we studied the description of $np$ scattering in the ${{}^{1}S}_{0}$ partial wave in the single-channel version of the Efros approach and in the ${{}^{3}SD}_{1}$ coupled partial waves in its two-channel version using the realistic Daejeon16 $NN$ interaction. To investigate the location of $S$-matrix poles, we explored the $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering described by the Woods–Saxon potential. We located also the $S$-matrix poles in the two-channel systems by calculating the deuteron bound state using the Daejeon16 $NN$ interaction and resonance parameters in the Noro–Taylor two-channel model Norro1980 .
Having in mind future applications in many-body systems in combination with the NCSM NCSMProg , we exploited the harmonic oscillator basis and introduced the Hamiltonian $H^{tr}$ truncated in the oscillator basis at some maximal oscillator quanta $N_{\max}$ as is done in the NCSM. We demonstrated how this truncated Hamiltonian $H^{tr}$ can be utilized within the Efros method. We suggested some modifications of the formalism of the Efros method using achievements developed within the HORSE formalism BangHORSE in scattering theory. The Efros method is equivalent to the HORSE approach if the set of SRFs includes all oscillator functions with oscillator quanta up to $N_{\max}$. This is, however, impractical in many-body applications since such a set of SRFs is extremely large. The great advantage of the Efros method is that it provides a reasonable convergence with significantly reduced SRF sets. One can use as SRFs a limited number of the oscillator functions describing the relative motion in each channel, a number of the eigenfunctions of $H^{tr\!}$, or their mixture. The use of the $H^{tr}$ eigenfunctions seems to be most convenient for future many-body applications of the Efros method in combination with the NCSM. Therefore, most of our studies were performed with the eigenfunction SRF sets. We demonstrated that the smoothing of potential energy matrix elements suggested in Refs. HungOne ; HungTwo significantly improves the convergence of all calculations.
Generally, the Efros method provides accelerated convergence and good accuracy in the description of various observables with a number of eigenfunction SRFs that are accessible in modern NCSM calculation. We did not find problems in calculations of single-channel scattering but in calculations of two-channel scattering, we obtained spikes and discontinuities of some $S$-matrix elements. These shortcomings, however, can be eliminated by reordering and rearranging the SRF set.
We conclude that the Efros method in combination with the NCSM is promising for ab initio studies of elastic scattering and reactions with light nuclei. By locating the $S$-matrix poles within this approach, one can investigate resonant states in light nuclei and derive resonant energies and widths. The $S$-matrix poles associated with bound states which can be calculated within the Efros extension of the NCSM, will improve the results of calculations of bound state energies and wave functions. In particular, it will be possible to describe the asymptotically decreasing tails of the wave functions. These improvements will result in converged and accurate predictions of various nuclear observables like the rms radii, quadrupole moments, asymptotic normalization constants, electromagnetic transition rates, etc. The most important improvements are expected in weakly-bound systems, e. g., in drip-line nuclei which are a focus of modern research in nuclear physics.
Acknowledgments
We are thankful to V. D. Efros for valuable discussions. This work was supported by US DOE Grants DE-SC0023692 and DE-SC0023707 under the Office of Nuclear Physics Quantum Horizons program for the “Nuclei and Hadrons with Quantum computers (NuHaQ)” project.

### v2

@@ III Two-channel scattering
In Fig. S19 the absolute magnitudes of the wave function expansion coefficients $d_{n,ij}(E)$ are shown at the energy of $E=0.5$ MeV where we obtain $\delta_{0}$, $\delta_{2}$, and $\epsilon$ close to exact value in calculations with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and $^{3}D_{1}$ as the first channel with $i=1$. We observe a fast convergence of the amplitudes even with small $N$ values. The arguments of the coefficients converge in a similar manner as seen in Fig. S20.
[…]
Figure S21 shows the convergence with $N_{\textrm{max}}$ of the resonance energy $E_{r}$ and width $\Gamma$ of $n\alpha$ scattering using the Woods–Saxon interaction with parameters fitted in Ref. WoodsSaxon1979 . The results are obtained by the HORSE method. We see that at $N_{\textrm{max}}=10$, we obtain results very close to the exact values.
Figure S21: $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering supported by the Woods–Saxon potential: a) resonance energy $E_{r}$ and b) resonance width $\Gamma$ obtained by locating the $S$-matrix pole within the HORSE method using various $N_{\max}$ truncations, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$.
@@
Figure S22 shows the convergence of the deuteron ground state energy $E_{b}$ with $N_{\textrm{max}}$ within the HORSE method. The deuteron ground state energy is seen to converge much faster to the exact value than the variational results obtained by the direct diagonalization of the truncated Hamiltonian $H^{tr}$. Approximately $N_{\textrm{max}}=50$ is needed to achieve three-digit accuracy for direct diagonalization as compared to $N_{\textrm{max}}=14$} needed for the same accuracy for the $S$-matrix pole location via the HORSE method.
@@
In Fig. S23, we present $\breve{d}_{n,i}(E_{b})$ of the deuteron wave function obtained by the adapted Efros method with different $N$ using eigenfunction SRFs, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$. We see a high accuracy of expansion coefficients calculated even with small $N$.
Figure S22: Deuteron bound state energy $E_{b}$ obtained with Daejeon16 $NN$ interaction using various $N_{\textrm{max}}$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ by the HORSE method and by the direct diagonalization of $H^{tr}$.
Figure S23: Absolute magnitudes of expansion coefficients $|\breve{d}_{n,i}(E_{b})|$ of the deuteron wave function obtained by locating the $S$-matrix pole within the adapted Efros method with different $N$ using eigenfunction SRFs, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$: a) $^{3}S_{1}$ components, b) $^{3}D_{1}$ components. The deuteron energy obtained with each NN is given in the legend in parentheses. We present also for comparison the exact values of $|\breve{d}_{n,i}(E_{b})|$ ($E_{b}=-2.225$ MeV) and those values obtained by the HORSE method. Symbols are often obscured by other symbols especially in panel (a).
Figure S24: a) Energy $E_{r}$ and b) width $\Gamma$ of the resonance in the two-channel Noro–Taylor model (see Ref. Norro1980 ) in HORSE calculations with $\Omega=1.5$ au, smoothing parameter $a=5$, and various $N_{\max}$.
@@
Figure S25: a) Energy $E_{r}$ and b) width $\Gamma$ of the resonance in the two-channel Noro–Taylor model (see Ref. Moiseyev1990 ) in calculations within the adapted Efros method with eigenfunction SRFs, $\Omega=1.5$ au, $N_{\max}=20$, smoothing parameter $a=5$, and various NN as compared with the exact results and with the HORSE method results.
$V_{ij}=\lambda_{ij}r^{2}e^{-r}+E_{i}^{\textrm{thresh.}}\delta_{ij},$ (S10a)
$\lambda_{11}=1,\quad\lambda_{12}=\lambda_{21}=-7.5,\quad\lambda_{22}=7.5,$ (S10b)
$E_{1}^{\textrm{thresh.}}=0,\quad E_{2}^{\textrm{thresh.}}=0.1.$ (S10c)
@@
The atomic units are used here. The interaction (S10) has a rich spectrum containing four bound states and a number of resonances, and is therefore a good test problem for the numerical validity of a particular approach. We explore the lowest-lying narrow resonance in this system at the energy $E_{r}=4.7682$} au with the width $\Gamma=0.00142$ au.
@@
Figure S24 shows the convergence with $N_{\textrm{max}}$ of the resonant energy $E_{r}$ and width $\Gamma$ in the Noro–Taylor two-channel model obtained by the HORSE method. $N_{\textrm{max}}=10$ is sufficient for obtaining $E_{r}$ with an accuracy of 1%. For the width $\Gamma$, approximately $N_{\textrm{max}}=50$ is needed to achieve a similar level of accuracy. Note, however, that the resonance is very narrow and the corresponding $\Gamma$ is small; thus the accuracy of 1% for the width means fixing it within a very small interval of absolute values. For both of these observables, we see a good level of stability beyond the truncation of $N_{\max}=22$.
@@
The results of calculations of $E_{r}$ and $\Gamma$ within the adapted Efros method with eigenfunction SRFs, $\Omega=1.5$ au, $N_{\max}=20$}, and smoothing parameter $a=5$ at various $N$ are presented in Fig. S25 and are compared with the HORSE method result. We see that the level of accuracy within better than 0.1% of the exact result is achieved for $E_{r}$ at $N=10$. For the width $\Gamma$, a relative accuracy of better than 15% is achieved at around $N=19$. We note that both the resonance energy $E_{r}$ and its width $\Gamma$ converge with $N$ to the HORSE results which require larger $N_{\max}$ values (see Fig. S24 ) to reproduce accurately the width. With $N=19$ we exactly reproduce the HORSE width; thus one needs to perform calculations with larger $N_{\max}$ to describe the $\Gamma$ value with higher accuracy.